(x/5) + (3/4) = 2x - 1
Answer by nerdybill(7384)

(x/5) + (3/4) = 2x - 1
.
Multiply both sides by 20 (common denominator):
4x + 15 = 40x - 20
Subtract 4x from both sides to get:
15 = 36x - 20
Add 20 to both sides to get:
35 = 36x
Finally, divide both sides by 36:
35/36 = x